Daytime

RRH installation

  • removed brackets and pins of the mirror cover. Stored in AO cabinet #2
  • installed brackets of RRH and installed RRH. No problems.

4D alignment

  • set HP and M3 at the position calibrated on Jun 23, 2010
  • shell set and flat using the default flattening of the last run: IRTC.400modes_20100623.sav.sav
  • The beam returned back to the F/15 focal plane just 2cm far from the interferometer focus
  • spot moved with Rx, Ry and z to superimpose the exit forcus with return focus. "By-eye" procedure
  • 4D shows fringes. WARNING: remind to check that 4D is working with FULL FRAME, fringes are dense in this stage of alignment and easily 4D provide wrong results with lower sampling because of aliasing
  • Used tool RR_at_telescope_v4.xls for computing hexapod commands to align the mirror from Zernike tip-tilt, focus and coma values provided by 4D
  • Final alignment provided:
    • Hexpaod:
      x y z Rx Ry Rz
      mm mm mm as as as
      0.662 3.051 3.370 -577.841 106.419 0.0
    • M3
      tip tilt pist sel.
      as as mm deg
      -612.1 739.8 0.0 8.5987
    • Mirror figure saved in "D:\4D\data\flattening\20100919\20100919_190000_first_alignment".
      • 4D-Zernike components:
        • AstigX: 3.0um (WFE)
        • AstigY: -3.2um
        • Focus: 0.70um
        • ComaX: -0.16um
        • ComaY: 0.04um
        • Higher orders: 280nm rms (wo trefoil: 180nm rms)

Flattening

  • the 4D does not be triggered
  • tested with spare fiber-to-BNC transceiver and related spare power supply (spare parts tested as functioning two days before in clean room with spare swing-arm electronics). NO trigger out of the transceiver (checked with scope)
  • test with a spare fiber from the LIU panel in the lower-right tree-house (the same of the transceiver) to the transceiver, NO trigger out of the transceiver
  • moved telescope horizon pointing and tested the trigger directly from the switch BCU: Trigger OK!
  • WARNING: the mirror has been left set by mistake running to horizon. Danger for the magnets!
  • THE PROBLEM IS THE FIBER: back to zenith, following the routing of the trigger fiber from the swing-arm cabinet we found the trigger fiber DISCONNECTED in the upper-left tree-house
  • procured a fiber for routing directly the trigger signal from the upper-left tree-house to the transceiver in the lower-right tree-house.
  • Test not done: 17:30 AZ time: time to dismount the RRH

RRH dismounting

  • smooth dismounting
  • brackets left on the HUB

Night

  • Swing arm retracted: no allowed to set the shell
  • dome opened: max wind speed 7m/s at 02:12UT

-- ArmandoRiccardi - 20 Sep 2010
Topic revision: r1 - 20 Sep 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback